OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution Those are the answers.

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

The first step is to isolate the absolute value expression.

OpenStudy (anonymous):

1. divide each side by -3, to get the abs(2x+6)=4 Now you let 4 be positive and negative. the new two equations are: 2x+6=4 and 2x+6=-4 Now solve for x which is -1 and -5.